Circular No. 7533

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2000ez IN NGC 3995 G. M. Hurst, Basingstoke, England, reports the discovery of an appar-

ent supernova (mag 16.8) by M. Armstrong, Rolvenden, on a CCD exposure obtained on Nov. 24.167 UT with a 0.35-m telescope in the course of the U.K. Nova/Supernova Patrol. SN 2000ez is located at α = 11^{h}57^{m}45^{s}.75, δ = +32^{o}17′49′′.1 (equinox 2000.0), which is 22′′.5 east and 12′′ north of the center of NGC 3995. The new object was confirmed on images taken by M. Schwartz (0.52-m telescope; mag 16.9) on Nov. 28.846, by P. Garnavich (Vatican Advanced Technology Telescope; R = 17.5) on Nov. 29.5, and by Armstrong on Nov. 30.731. Garnavich adds that SN 2000ez appears red and may be extinguished by dust in the host galaxy. SN 2000ez does not appear on second Palomar Sky Survey images taken on 1991 Apr. 3 (red plate; limiting mag 20.8) and on 1996 Mar. 17 (blue plate; limiting mag 22.5), or on CCD images obtained by T. Boles (1999 Dec. 5) and by the Lick Observatory Supernova Search (1999 May 8).

SUPERNOVA 2000fa IN UGC 3770 A. Friedman and W. D. Li, University of California at Berkeley, on be-

half of LOTOSS (cf. IAUC 7514), report the discovery with the 0.8-m Katzman Automatic Imaging Telescope (KAIT) of an apparent supernova (mag ∼ 17.5) on an unfiltered image taken on Nov. 30.5 UT. SN 2000fa was confirmed (at similar brightness) on an unfiltered image taken by M. Schwartz with the 0.5-m Tenagra III automatic telescope on Dec. 1.2. The new object is located at α = 7^{h}15^{m}29^{s}.79, δ = +23^{o}25′32′′.5 (equinox 2000.0), which is 5′′.1 east and 5′′.6 south of the nucleus of UGC 3770. KAIT images of the same field taken on Nov. 24.5 (limiting mag ∼ 18) and 17.4 (limiting mag ∼ 19.0) showed nothing at this position.

SUPERNOVA 2000el IN NGC 7290 T. Matheson, S. Jha, P. Challis, and R. Kirshner, Harvard-Smithsonian

Center for Astrophysics, report that a spectrum of SN 2000el (cf. IAUC 7523, 7531), obtained by P. Berlind on Nov. 25.13 UT with the F. L. Whipple Observatory 1.5-m telescope (+ FAST spectrograph), shows it to be a type-II supernova several days after maximum light. The spectrum consists of a well-defined continuum with superposed P-Cyg lines of hydrogen and Ca II. Narrow Hα emission superposed on the spectrum of the supernova indicates a recession velocity of 2970 km/s for the host galaxy. The expansion velocity derived from the minimum of the Hβ line is 4200 km/s.

2000 December 1  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green